A researcher wishes to estimate with 99% confidence, the population proportion of adults who eat fast food four to six times per week.Her estimate must be accurate within 5% of the population proportion.
a) no preliminary estimate is avaiable, find the minimum sample size needed.
b) find the minimium sample size needed, using a prior study that found that 38% of the respondents said they eat fast food four to six times per week
c) compare the results from parts (a) and (b)
Answer #1

Sample size = Z2/2 * p ( 1 - p) / E2

a)

When prior estimate for proportion is not specified then p = 0.50

Sample size = 2.57582 * 0.5 * 0.5 / 0.052

= 663.5

n = 664 (Rounded up to nearest integer)

b)

Sample size = 2.57582 * 0.38 * (1 - 0.38) / 0.052

= 625.26

n = 626 (Rounded up to nearest integer)

c)

With known estimate of proportion in part b) dramatically reduces the sample size.